OpenStudy (anonymous):

Solve the equation for y: 4x + 3y = -12

OpenStudy (anonymous):

y=-4-4/3x

hero (hero):

4x + 3y = -12 3y = -12 - 4x 3y = -(12 + 4x) y = -(12 + 4x)/3 y = -4 - 4x/3

OpenStudy (anonymous):

slope intercept form ( y=mx+b )

hero (hero):

second option